OpenStudy (anonymous):

The length of a diagonal of a square is 5 inches. Find the perimeter of the square.

OpenStudy (anonymous):

Let d be the length of every side of a square. then the perimeter = 4*d inches Using Pythagoras Theorem, x^2+y^2 = z^2 where x=y=d and z = diagonal length = 5 d^2+d^2=5^2 2*d^2 = 25 d^2 = 25/2 = 12.5 so d = sqrt(12.5) = 3.5335... so perimeter = 4*sqrt(12.5)= 14.14